OpenStudy (anonymous):

Please help me understand the problem 2 + 4 * 3^x =4, solve for x ?

OpenStudy (psymon):

\[2+4(3^{x})= 4\] Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I see 4(3^x) = 2, but I don't know how to bring down x, log perhaps???

OpenStudy (psymon):

Get the exponential part by itself as much as possible first. So now you would divide by 4 to get 3^x by itself.

OpenStudy (anonymous):

ok! thats an idea 3^x = 2/3

OpenStudy (psymon):

Should be (1/2) on the otherside xD

OpenStudy (anonymous):

ugh you are right Psymom, so can I do x log 3 = log 1/2 ???

OpenStudy (anonymous):

*Psymon

OpenStudy (psymon):

Well, it would have to be calculator work. So yeah, you'd divide both sides by ln3

OpenStudy (anonymous):

x = (log1/2)log 3 ??? I'm running the calc now

OpenStudy (anonymous):

If a take log on one side, don't I also apply that to the other side?

OpenStudy (psymon):

Well, you started with this before the ln: \[3^{x} = \frac{ 1 }{ 2 }\] Then we took the natural log of both sides to help get the x out of the exponent position: \[xln(3) = \ln(\frac{ 1 }{ 2 })\] Now I would divide both sides by ln(3) to get: \[x = \frac{ \ln(\frac{ 1 }{ 2 }) }{ \ln(3) }\approx -.691\]

OpenStudy (isaiah.feynman):

I hate myself

OpenStudy (anonymous):

you both have been helpful! We're all here to learn and help. Thank you both and thanks Psymon, I was getting the idea but my computation was wrong. Thanks for the correction!

OpenStudy (psymon):

Yep, np ^_^

OpenStudy (anonymous):

:-)

OpenStudy (isaiah.feynman):

Giving an erroneous calculation is not helpful!